OpenStudy (lxelle):

Compressed air is escaping from a container. The pressure of the air in the container at time t is P, and the constant atmospheric pressure of the air outside the container is A. The rate of decrease of P is proportional to the square root of the pressure difference (P-A). Thus the differential equation connecting P and t is dP/dt = -k sqrt(P-A), where k is a positive constant.

OpenStudy (lxelle):

i) Find, in any form, the general solution of this differential equation. (ii) Given that P = 5A when t = 0, and that P = 2A when t = 2, show that k = sqrtA. (iii) Find the value of t when P = A. (iv) Obtain an expression for P in terms of A and t.

OpenStudy (sleepyjess):

I'm not sure, but @Kainui is pretty awesome

OpenStudy (kainui):

Yeah I can help you with this, let's start with whichever part you want first, show me what you can do or what you know already.

OpenStudy (lxelle):

i)

OpenStudy (lxelle):

I dont get what do they mean by the general solution

OpenStudy (unklerhaukus):

solve the DE; its variables are seperable

OpenStudy (kainui):

General solution just means solve the differential equation but without solving for the constants of integration. For example, if I had the differential equation \[\frac{dy}{dx} = 2x+5\] then the general solution would be \[\Large y = x^2+5x+C\]

OpenStudy (lxelle):

i dont know how to work that dP/dt out. D:

OpenStudy (unklerhaukus):

\[\quad\qquad \frac{\mathrm dP}{\mathrm dt} = -k \sqrt{P-A}\\ \quad\frac{\mathrm dP}{\sqrt{P-A}} = -k\,\mathrm dt\\ \int\frac{\mathrm dP}{\sqrt{P-A}} = \int-k\,\mathrm dt\\ \qquad\qquad\quad =\]

OpenStudy (unklerhaukus):

perform the integrations, the one on the left can be found using a table of standard integrals, the one on the right is simple

OpenStudy (lxelle):

I got 1/ sqrt (P-A), i got the rhs. but dont really know how to continue lhs. D:

OpenStudy (lxelle):

@UnkleRhaukus

OpenStudy (unklerhaukus):

For the left hand side use:\[\int\frac{\mathrm dx}{\sqrt{x\pm a}}=2\sqrt{x\pm a}+c_1\] For the right hand side use:\[\int a\,\mathrm dx = a\int\mathrm dx = ax+c_2\]

OpenStudy (lxelle):

I dont get how the 2 appears there

OpenStudy (lxelle):

Anyways help me with ii? @UnkleRhaukus

OpenStudy (lxelle):

I got 2 sqrtA = -k + 2 sqrt4A What should i do next

OpenStudy (perl):

$$ \large \int\frac{\mathrm dx}{\sqrt{x\pm a}} \\ \large = \int (x \pm a)^{-1/2}~dx \\ \large = \frac{(x \pm a)^{1/2}}{\frac{1}{2}} = 2(x \pm a)^{1/2} $$

OpenStudy (lxelle):

Nice @perl thanks.

OpenStudy (unklerhaukus):

don't forget the constant of integration

OpenStudy (perl):

yes :) sorry

OpenStudy (lxelle):

I'm done with part i. What about the ii

OpenStudy (unklerhaukus):

What did you get for part i) ?

OpenStudy (unklerhaukus):

[(it should have a \(t\) in it]

OpenStudy (lxelle):

2 SQRT (P-A) = -kt+c

OpenStudy (unklerhaukus):

that's good

OpenStudy (lxelle):

i work it out until 2 srqt A = -2k + 2sqrt4A then stuck

OpenStudy (unklerhaukus):

Your general solution is: \[2 \sqrt{P-A} = -kt+c\] The initial condition is that: \[P = 5A\quad\text{ when }\quad t = 0\]sub this into the general solution, to solve for \(c\)

OpenStudy (lxelle):

i got 2 sqrt 4A for c

OpenStudy (unklerhaukus):

yes \[2\sqrt{4A} = c\] but you can simplify this

OpenStudy (lxelle):

simplify to?

OpenStudy (lxelle):

4sqrtA?

OpenStudy (unklerhaukus):

yep, so \(c=4\sqrt A\) plugging this into the general solution, we have a particular solution: \[2 \sqrt{P-A} = -kt+4\sqrt A\]

OpenStudy (lxelle):

I did alrd. then?

OpenStudy (unklerhaukus):

We also know that \(P = 2A\) when \(t = 2\), plug these into the particular solution to find the constant of proportionality \(k\),

OpenStudy (lxelle):

yeah yeah i did lol.

OpenStudy (lxelle):

2 srqt A = -2k + 4sqrrt A

OpenStudy (unklerhaukus):

now solve for \(k\) . . .

OpenStudy (lxelle):

THANKSS.

OpenStudy (lxelle):

@UnkleRhaukus I need help with iv!

OpenStudy (unklerhaukus):

what did you get for part iii)

OpenStudy (lxelle):

t = 4

OpenStudy (unklerhaukus):

and did you get k=sqrt A for ii) ?

OpenStudy (lxelle):

I substitute them in, i got 2 sqrt(P-A) = -sqrt A t+4sqrt A

OpenStudy (unklerhaukus):

\[2 \sqrt{P-A} = -\sqrt A t+4\sqrt A\] good, now, can you solve this for \(P\)?

OpenStudy (lxelle):

No. D:

OpenStudy (unklerhaukus):

factor the right hand side first,

OpenStudy (unklerhaukus):

then divide by 2, and square both sides

OpenStudy (lxelle):

i can't get the answer. D:

OpenStudy (unklerhaukus):

what do you get when you factor \(-\sqrt At+4\sqrt A\) ?

OpenStudy (lxelle):

Sqrt A (4- t)

OpenStudy (lxelle):

ooh got it!! thanks.

OpenStudy (unklerhaukus):

so what is \(P\) in terms of \(A\) and \(t\) ?

OpenStudy (lxelle):

1/4 A (4+(4-t)^2)

OpenStudy (unklerhaukus):

good work!

OpenStudy (lxelle):

:)